OpenStudy (anonymous):

split into partial fractions: 4/(y^3-25y)

OpenStudy (anonymous):

-4/(25 y)+2/(25 (y+5))+2/(25 (y-5))

OpenStudy (cwrw238):

4/(y^3-25y) = A/y + B/y-5 + C/ y+5 4 = A((y-5)(y+5) + By(y+5) + Cy(y-5) let y = 5 4 = 50B B = 2/25 let y = -5 4 = 50C C = 2/25 let y = 0 4 = -25A A = -4/25 so its -4 2 2 ---------- + ---------- + ------- 25y 25(y+5) 25(y-5) if my algebra is correct

OpenStudy (cwrw238):

the way to check is to work back and see if you get the original exprersion

OpenStudy (cwrw238):

- oh laurly got same answer